Returns.

What are the returns on the following​ investments?  

Investment

Original Cost

of Investment

Selling Price

of Investment

Distributions

Received

Percent Return

  CD

​ $600

​$650

​$0

​?

  Stock

​$34

​$43

​$33

​?

  Bond

1060

​1010

​$60

​?

  Bicycle

​$310

​$210

​$0

​?
Answer #1

CD =650/600-1 = 8.33%

Stock =(43+33)/34-1 =123.53%

Bond =(1010+60)/1060-1 =0.94%

Bicycle =210/310-1 =-32.26%